OpenStudy (anonymous):

A mass suspended on a spring will exhibit sinusoidal motion when it moves. If the mass on a spring is 85 cm off the ground at its highest position and 41 cm off the ground at its lowest position and takes 3.0 s to go from the top to the bottom and back again, determine an equation to model the data.

OpenStudy (anonymous):

\[f(x)=44\sin (2\pi/3)\]

OpenStudy (anonymous):

jk... \[f(x)=22\cos(2\pi/3)+63\]